# Online Conformal Probabilistic Numerics via Adaptive Edge-Cloud Offloading

### v1

@@ Abstract
Consider an edge computing setting in which a user submits queries for the solution of a linear system to an edge processor, which is subject to time-varying computing availability. The edge processor applies a probabilistic linear solver (PLS) so as to be able to respond to the user’s query within the allotted time and computing budget. Feedback to the user is in the form of an uncertainty set. Due to model misspecification, the uncertainty set obtained via a direct application of PLS does not come with coverage guarantees with respect to the true solution of the linear system. This work introduces a new method to calibrate the uncertainty sets produced by PLS with the aim of guaranteeing long-term coverage requirements. The proposed method, referred to as online conformal prediction-PLS (OCP-PLS), assumes sporadic feedback from cloud to edge. This enables the online calibration of uncertainty thresholds via online conformal prediction (OCP), an online optimization method previously studied in the context of prediction models. The validity of OCP-PLS is verified via experiments that bring insights into trade-offs between coverage, prediction set size, and cloud usage.
@@ 1 Introduction
In modern hierarchical computing architectures encompassing edge and cloud segments, the efficient management of computational resources is critical to balance performance with latency and communication overhead. For instance, in industrial Internet-of-Things (IoT) settings, sensor data processing must integrate real-time decision-making at the edge with more complex analytics in the cloud, while accounting for inherent trade-offs between response speed and accuracy (Hong et al., 2019).
Linear systems serve as the cornerstone of virtually all numerical computation. These systems – in the form of the equation $Ax=b$ with unknown vector $x$ – arise in contexts including convex optimization (Boyd and Vandenberghe, 2004), Kalman filtering for state estimation (Thrun, 2002), and finite element analysis for computational fluid dynamics (Gassner and Winters, 2021). A timely and efficient solution of such systems is often critical for real-time decision-making.
Figure 1: At each round $t$, a user submits a linear system defined by the pair $(A_{t},b_{t})$ to an edge device. Given the available computing budget $I_{t}$, the edge device employs a probabilistic linear solver (PLS), obtaining an uncertainty set $\mathcal{C}_{t}$ for the true solution $x^{*}_{t}=A_{t}^{-1}b_{t}$. The set $\mathcal{C}_{t}$ is returned in a timely fashion to the user. The proposed method, OCP-PLS ensures long-term coverage guarantees (2) for the uncertainty sets $\mathcal{C}_{t}$, addressing model misspecification in PLS. To this end, OCP-PLS allows for sporadic communication between cloud and edge.
@@
Iterative solvers provide an ideal solution to adapt to the available computing budget, as they refine the solution sequentially until the computing budget runs out. Specifically, probabilistic numerics (PN) (Larkin, 1972) treats numerical problems, such as solving linear systems, as a form of statistical inference. This provides a natural way to quantify the uncertainty associated with iterative numerical solutions obtained under limited computing power.
Specifically, probabilistic linear solvers (PLSs) (Cockayne et al., 2019a; Hennig et al., 2015; Bartels et al., 2019; Reid et al., 2020; Pförtner et al., 2024) treat the solution of a linear system as a Bayesian inference problem, whereby each iteration provides an updated posterior distribution for the solution $x^{*}_{t}$. The posterior distribution reflects the current knowledge of the true solution given the available computing budget. Using this posterior distribution, PLS can construct an uncertainty set $\mathcal{C}_{t}$ with the aim of covering the true solution $x^{*}_{t}$ with a pre-determined target probability level.
In the considered system illustrated in Fig. 1, at each round $t$, the edge processor applies PLS so as to be able to respond to the user’s query within the allotted time and computing budgets. Feedback to the user is in the form of an uncertainty set $\mathcal{C}_{t}$. However, in the presence of model misspecification, the uncertainty sets obtained using a direct application of PLS do not come with coverage guarantees with respect to the true solution $x^{*}_{t}$ of the linear system (Cockayne et al., 2019a; Bartels et al., 2019; Wenger and Hennig, 2020; Hennig et al., 2015), and are often too conservative (Hegde et al., 2024; Reid et al., 2020; Wenger and Hennig, 2020). Note that misspecification in PLS often occurs during its iterative update procedure, which comes as a cost to accelerate the computing speed (Cockayne et al., 2019a, 2022, b).
This work introduces a new method to calibrate the uncertainty sets produced by PLS with the aim of guaranteeing long-term coverage requirements, while addressing the excessive conservativeness of PLS. In practice, we wish to ensure that, on average over time, the uncertainty set $\mathcal{C}_{t}$ returned by the edge to the user contains the true solution $x^{*}_{t}$ with a user-defined coverage rate $1-\alpha$.
@@ 2.1 Setting and Problem Definition
As illustrated in Fig. 1, we study a hierarchical computing architecture encompassing edge and cloud processors. At each round $t$, a user submits a linear system problem defined by the pair $(A_{t},b_{t})$, consisting of an $n_{t}\times n_{t}$ Hermitian matrix $A_{t}$ and an $n_{t}\times 1$ vector $b_{t}$, to the edge device. The user is interested in obtaining information about the optimal solution
@@
At each round $t$, the edge has available a computing budget $I_{t}$. Accounting for this limited budget, the edge device employs PLS to obtain an uncertainty set $\mathcal{C}_{t}$ for the true solution (1) (Cockayne et al., 2019a; Bartels et al., 2019; Wenger and Hennig, 2020). This set is returned to the user in response to its query. Thanks to processing at the edge, the uses’ requests experience low latency (see Fig. 2).
This paper addresses the problem of calibrating the prediction set $C_{t}$ so as to ensure that the sequence of uncertainty sets $C_{t}$ returned to the user satisfies the following long-term coverage condition
$\left|\frac{1}{T}\sum_{t=1}^{T}\Pr\{x_{t}^{*}\in\mathcal{C}_{t}\}-(1-\alpha)\right|\leq o(1),$ (2)
@@ 2.2 Probabilistic Linear Solver
Given a budget of $I_{t}$ iteration, PLS operates sequentially by considering at each iteration $i=1,\ldots,I_{t}$ a search direction $s_{i}$, yielding an effective observation $y_{i}=s_{i}^{T}b$. Given a Gaussian prior distribution $x^{*}\sim\mathcal{N}(\mu_{0},\Sigma_{0})$ on the solution $x^{*}=A_{t}^{-1}b_{t}$, PLS returns the Gaussian posterior
@@
collecting the $I_{t}$ search directions.
Given the posterior distribution (3), an $(1-\alpha)$-uncertainty set for the true solution $x^{*}_{t}$ can be obtained as
@@
The uncertainty set (2.2) covers the true solution $x^{*}_{t}$ in (1) with probability no smaller than $1-\alpha$, i.e.,
@@
only if the posterior distribution $p(x_{t}|A_{t},b_{t})$ is able to describe the true uncertainty on the solution $x^{*}_{t}$ given the observation $y_{1},\dots,y_{I_{t}}$, i.e., $p(x_{t}|A_{t},b_{t})$ is well calibrated.
@@ 3.1 Online Conformal Prediction
Given an arbitrary sequence of input-output pairs $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$, for $t=1,2,\ldots$, OCP aims to construct a prediction set $\mathcal{C}_{t}(x_{t})$ on the output space $\mathcal{Y}$ that satisfies long-term coverage guarantees. OCP builds on a scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ that measures the extent to which the output $y$ is mismatched to the input $x$. A smaller value $s(x,y)$ indicates that $y$ is a better fit for input $x$. The set $\mathcal{C}_{t}(x_{t})$ is obtained by choosing all the candidate output values $y\in\mathcal{Y}$ with sufficiently small scores $s(x,y)$, i.e.,
@@ 3.2 Intermittent Online Conformal Prediction
The OCP update (10) requires dense feedback, as it assumes the availability of the true output $y_{t}$ at every time step $t=1,2,...$ in (10). Intermittent OCP (I-OCP) (Zhao et al., 2024) alleviates this assumption by leveraging inverse probability weighting, also known as the Horvitz-Thompson estimator, which is used in statistical inference with missing data (Imbens and Rubin, 2015).
@@
It can be proved that I-OCP satisfies the long-term expected coverage property
@@ 4 Online Conformal Probabilistic Linear Solver
In this section, we present OCP-PLS, a novel framework for edge inference via PN that satisfies the quality-of-service requirement (2) by integrating PLS (see in Section 2.2) with I-OCP (see Section 3).
@@ 4.1 Score Function
As discussed in Section 3.1, OCP requires the definition of a score function for any value of the target variable. In the setting described in Section 2, the target variable at each time $t$ is the true solution $x^{*}_{t}$ in (1). Given the posterior distribution (3)-(4) produced by PLS, we set the score for any candidate solution $x_{t}$ to be proportional to the density $\mathcal{N}(x_{t}|\mu_{t},\Sigma_{t})$ as
$s(x_{t};A_{t},b_{t})=\text{exp}\left(-(x_{t}-\mu_{t})^{T}\Sigma_{t}^{-1}(x-\mu_{t})\right).$ (19)
Note that the score (19) depends on the specific sequence of search directions (5), although we do not make this dependence explicit in the notation $s(x_{t};A_{t},b_{t})$.
@@
$\mathcal{C}_{t}=\{x_{t}\in\mathbb{R}^{d}:s(x_{t}:A_{t},b_{t})\geq\lambda_{t}\},$ (20)
@@
Second, computing the inverse $\Sigma_{t}^{-1}$ in (19) generally requires a complexity of the order $O(n_{t}^{3})$, which would not be feasible under the edge budget constraint. However, given the structure of the covariance matrix $\Sigma_{t}$ in (4) and assuming the use of conjugate gradient search directions (Cockayne et al., 2019a; Bartels et al., 2019), the inverse $\Sigma_{t}^{-1}$ can be computed efficiently using the Sherman-Morrison-Woodbury formula (Miller, 1981) at a complexity of the order $\mathcal{O}(I_{t}n_{t}^{2})$ (Hegde et al., 2024).
@@ 4.2 Cloud-to-Edge Feedback and Threshold Selection
As illustrated in Fig. 1 and Fig. 2, in order to set the threshold $\lambda_{t}$ in (20), OCP-PLS leverages sporadic feedback from the cloud to the edge. Specifically, we propose to submit a query $(A_{t},b_{t})$ to the cloud with a probability $p_{t}$ that increases with the size of the uncertainty set $\mathcal{C}_{t}$, i.e.,
$|\mathcal{C}_{t}|=\frac{\pi^{n_{t}/2}}{\Gamma(n_{t}/2+1)}\cdot(\text{det}(\Sigma_{t}))^{1/2}\cdot(-\log\lambda_{t})^{n_{t}/2},$ (21)
where $\Gamma(\cdot)$ is the Gamma function and $\text{det}(\cdot)$ is the determinant operator. In this way, cloud-to-edge feedback is more likely to be requested at time instants $t$ characterized by a larger uncertainty, hence making cloud feedback more valuable for calibration. Conversely, when the uncertainty set is small enough, we reduce the probability of requesting feedback, thereby saving cloud computing resources.
@@
$p_{t}=\sigma\left(\frac{\log|\mathcal{C}_{t}|}{n_{t}}-\theta\right),$ (22)
which is a function of the log value of the set size normalized by the matrix dimension $n_{t}$. Lower values of hyperparameter $\theta$ in (22) increase the probability of cloud-to-edge feedback, yielding a smaller set $\mathcal{C}_{t}$ at the cost of higher usage of cloud resources.
@@
10:     else
11:         $\lambda_{t+1}=\lambda_{t}$
12:     end if
13:end for

### v2

@@ Abstract
Consider an edge computing setting in which a user submits queries for the solution of a linear system to an edge processor, which is subject to time-varying computing availability. The edge processor applies a probabilistic linear solver (PLS) so as to be able to respond to the user’s query within the allotted time and computing budget. Feedback to the user is in the form of a set of plausible solutions. Due to model misspecification, the highest-probability-density (HPD) set obtained via a direct application of PLS does not come with coverage guarantees with respect to the true solution of the linear system. This work introduces a new method to calibrate the HPD sets produced by PLS with the aim of guaranteeing long-term coverage requirements. The proposed method, referred to as online conformal prediction-PLS (OCP-PLS), assumes sporadic feedback from cloud to edge. This enables the online calibration of uncertainty thresholds via online conformal prediction (OCP), an online optimization method previously studied in the context of prediction models. The validity of OCP-PLS is verified via experiments that bring insights into trade-offs between coverage, prediction set size, and cloud usage.
@@ 1 Introduction
In modern hierarchical computing architectures encompassing edge and cloud servers, the efficient management of computational resources is critical to balance performance with latency and communication overhead. Edge servers process data near the user, providing low-latency, but possibly inaccurate, responses due to limited computational resources. In contrast, cloud servers offer more powerful processing capabilities, but at the cost of introducing communication delays. As an example, in industrial Internet-of-Things (IoT) settings, sensor data processing may leverage both real-time decision-making at the edge and more complex data analytics in the cloud, while accounting for the resulting inherent trade-offs between response speed and accuracy (Hong et al., 2019).
Linear systems serve as the cornerstone of virtually all numerical computation. These systems – in the form of the equation $Ax=b$ with unknown vector $x$ – arise in contexts including convex optimization (Boyd and Vandenberghe, 2004), Kalman filtering for state estimation (Thrun, 2002), and finite element analysis for computational fluid dynamics (Gassner and Winters, 2021). A timely and efficient solution to such systems is often critical for real-time decision-making.
@@
Iterative solvers provide an ideal solution to adapt to the available computing budget, as they refine the solution sequentially until the computing budget runs out. Probabilistic numerics (PN) (Larkin, 1972) treats numerical problems, such as solving linear systems, as a form of statistical inference. This provides a natural way to quantify the uncertainty associated with iterative numerical solutions obtained under limited computing power.
Specifically, the probabilistic linear solver (PLS) (Cockayne et al., 2019a; Hennig et al., 2015; Bartels et al., 2019; Reid et al., 2020; Pförtner et al., 2024) typically treat the solution of a linear system as a Bayesian inference problem, whereby each iteration provides an updated posterior distribution for the solution $x^{*}_{t}$. The posterior distribution reflects the current knowledge of the true solution given the available computing budget. Using this posterior distribution, PLS can construct a highest-probability-density (HPD) set $\mathcal{C}_{t}$ with the aim of covering the true solution $x^{*}_{t}$ with a pre-determined target probability level.
In the considered system illustrated in Fig. 1, at each round $t$, the edge processor applies PLS so as to be able to respond to the user’s query within the allotted time and computing budgets. Feedback to the user is in the form of a HPD set $\mathcal{C}_{t}$. However, in the presence of model misspecification, the HPD sets obtained using a direct application of PLS do not come with coverage guarantees with respect to the true solution $x^{*}_{t}$ of the linear system (Cockayne et al., 2019a; Bartels et al., 2019; Wenger and Hennig, 2020; Hennig et al., 2015).
Figure 1: At each round $t$, a user submits a linear system defined by the pair $(A_{t},b_{t})$ to an edge device. Given the available computing budget $I_{t}$, the edge device employs a probabilistic linear solver (PLS), obtaining a highest-probability-density (HPD) set $\mathcal{C}_{t}$ for the true solution $x^{*}_{t}=A_{t}^{-1}b_{t}$. The set $\mathcal{C}_{t}$ is returned in a timely fashion to the user. The proposed method, OCP-PLS ensures long-term coverage guarantees (2) for the HPD sets $\mathcal{C}_{t}$, addressing model misspecification in PLS. To this end, OCP-PLS allows for sporadic communication between cloud and edge.
Model misspecification arises from the choice of a distribution that does not reflect the properties of the true solution $x^{*}_{t}$. For example, one typically assumes a Gaussian prior, but the solution $x^{*}_{t}$ may be sparse. This problem is exacerbated if PLS is implemented with prior-dependent search directions, such as BayesCG
(Hegde et al., 2024; Reid et al., 2020; Wenger and Hennig, 2020; Cockayne et al., 2019a, 2022, b), which may reinforce incorrect prior assumptions. Further work on the impact of calibration on PLS includes reference (Reid et al., 2020), which demonstrated that BayesCG search directions yield a slightly optimistic HPD set under the Krylov prior, and references (Hegde et al., 2024; cockayne2021probabilistic), which explored the use of probabilistic stationary iterative methods rather than Bayesian framework.
This work introduces a new method to calibrate the HPD sets produced by PLS with the aim of guaranteeing long-term coverage requirements. In practice, we wish to ensure that, on average over time, the HPD set $\mathcal{C}_{t}$ returned by the edge to the user contains the true solution $x^{*}_{t}$ with a user-defined coverage rate $1-\alpha$.
@@ 2.1 Setting and Problem Definition
As illustrated in Fig. 1, we study a hierarchical computing architecture encompassing edge and cloud processors. At each round $t$, a user submits a linear system problem defined by the pair $(A_{t},b_{t})$, consisting of an $n_{t}\times n_{t}$ Hermitian matrix $A_{t}$ and an $n_{t}\times 1$ vector $b_{t}$, to the edge device. The user is interested in obtaining information about the solution
@@
At each round $t$, the edge has available a computing budget $I_{t}$. Accounting for this limited budget, the edge device employs PLS to obtain an HPD set $\mathcal{C}_{t}$ for the true solution (1) (Cockayne et al., 2019a; Bartels et al., 2019; Wenger and Hennig, 2020). This set is returned to the user in response to its query. Thanks to processing at the edge, the user’s requests experience low latency (see Fig. 2).
This paper addresses the problem of calibrating the prediction set $\mathcal{C}_{t}$ so as to ensure that the sequence of HPD sets $\mathcal{C}_{t}$ returned to the user satisfies the following long-term coverage condition
$\left|\frac{1}{T}\sum_{t=1}^{T}\Pr\{x_{t}^{*}\in\mathcal{C}_{t}\}-(1-\alpha)\right|=o(1),$ (2)
@@ 2.2 Probabilistic Linear Solver
Given a budget of $I_{t}$ iterations, PLS operates sequentially by considering at each iteration $i=1,\ldots,I_{t}$ a search direction $s_{i}$, yielding an effective observation $y_{i}=s_{i}^{T}b$. Given a Gaussian prior distribution $x^{*}\sim\mathcal{N}(\mu_{0},\Sigma_{0})$ on the solution $x^{*}=A_{t}^{-1}b_{t}$ with a full-rank diagonal matrix $\Sigma_{0}$, PLS returns the Gaussian posterior
@@
collecting the $I_{t}$ search directions. Note that the covariance matrix $\Sigma_{t}$ has rank
$r_{t}\leq n_{t}-I_{t},$ (6)
with equality achieved if the directions $S_{t}$ are suitably chosen (Wenger et al., 2022). Note also that, due to the propagation of floating point error, the rank $r_{t}$ may be practically larger than $n_{t}-I_{t}$. This may be the case, e.g., when using BayesCG as the search directions and the number of iterations $I_{t}$ is close to the dimension $n_{t}$ (Cockayne et al., 2019a).
Given the posterior distribution (3), an $(1-\alpha)$-HPD set for the true solution $x^{*}_{t}$ can be obtained as
@@
The HPD set (2.2) covers the true solution $x^{*}_{t}$ in (1) with probability no smaller than $1-\alpha$, i.e.,
@@
only if the posterior distribution $p(x_{t}|A_{t},b_{t})$ accurately quantifies the uncertainty associated with estimating the solution $x^{*}_{t}$ given the observation $y_{1},\dots,y_{I_{t}}$. Accordingly, the posterior distribution $p(x_{t}|A_{t},b_{t})$ is regarded as well calibrated when the coverage probability in (8) equals the target level $1-\alpha$.
@@ 3.1 Online Conformal Prediction
Given an arbitrary sequence of input-output pairs $(x_{t},y_{t})\\ \in\mathcal{X}\times\mathcal{Y}$, for $t=1,2,\ldots$, OCP aims to construct a prediction set $\mathcal{C}_{t}(x_{t})$ on the output space $\mathcal{Y}$ that satisfies long-term coverage guarantees. OCP builds on a scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ that measures the extent to which the output $y$ is mismatched to the input $x$. A smaller value $s(x,y)$ indicates that $y$ is a better fit for input $x$. The set $\mathcal{C}_{t}(x_{t})$ is obtained by choosing all the candidate output values $y\in\mathcal{Y}$ with sufficiently small scores $s(x,y)$, i.e.,
@@ 3.2 Intermittent Online Conformal Prediction
The OCP update (11) requires dense feedback, as it assumes the availability of the true output $y_{t}$ at every time step $t=1,2,...$ in (11). Intermittent OCP (I-OCP) (Zhao et al., 2024) alleviates this assumption by leveraging inverse probability weighting, also known as the Horvitz–Thompson estimator, which is used in statistical inference with missing data (Imbens and Rubin, 2015).
@@
It can be proved that I-OCP satisfies the long-term expected coverage property (Zhao et al., 2024)
@@ 4 Online Conformal Probabilistic Linear Solver
In this section, we present OCP-PLS, a novel framework for edge inference via PN that satisfies the long-term quality-of-service requirement (2) by integrating PLS (see in Section 2.2) with I-OCP (see Section 3).
@@ 4.1 Score Function
As discussed in Section 3.1, OCP requires the definition of a score function for any value of the target variable. In the setting described in Section 2, the target variable at each time $t$ is the true solution $x^{*}_{t}$ in (1). Given the posterior distribution (3)-(4) produced by PLS, we set the score for any candidate solution $x_{t}$ to be proportional to the density of the scaled degenerate multivariate Gaussian distribution $\mathcal{N}(x_{t}|\mu_{t},\Sigma_{t})$. Let $\Sigma_{t}=V_{t}\Lambda_{t}V_{t}^{T}$ be the eigendecomposition of the rank-$r_{t}$ covariance matrix $\Sigma_{t}$, where $V_{t}$ is a unitary $n_{t}\times r_{t}$ matrix and $\Lambda_{t}$ is an $r_{t}\times r_{t}$ diagonal matrix. The score can be explicitly written as
$s(x_{t};A_{t},b_{t})=\left\{\begin{array}[]{ll}\exp(-(x_{t}-\mu_{t})^{T}V_{t}\Lambda_{t}^{-1}V_{t}^{T}(x_{t}-\mu_{t}))&\text{if }(x_{t}-\mu_{t})\in\text{range}(\Sigma_{t}),\\[10.00002pt] 0&\text{otherwise}.\end{array}\right.$
By doing this, the score function is bounded between 00 and $1$.
Since matrix $\Sigma_{t}$ in (4) can be expressed as a sequence of rank-$1$ updates to the prior diagonal matrix $\Sigma_{0}$, its eigendecomposition can be efficiently updated using
Given’s rotations (golub2013matrix, Section 12.5). This results in a per-iteration complexity of $\mathcal{O}(n_{t}^{2})$, and an overall complexity of $\mathcal{O}(I_{t}n_{t}^{2})$ after $I_{t}$ updates. Note that the score (20) depends on the specific sequence of search directions (5), although we do not make this dependence explicit in the notation $s(x_{t};A_{t},b_{t})$.
@@
$\mathcal{C}_{t}=\{x_{t}\in\mathbb{R}^{d}:s(x_{t};A_{t},b_{t})\geq\lambda_{t}\},$ (21)
@@ 4.2 Cloud-to-Edge Feedback and Threshold Selection
As illustrated in Fig. 1 and Fig. 2, in order to set the threshold $\lambda_{t}$ in (21), OCP-PLS leverages sporadic feedback from the cloud to the edge. Specifically, we propose to submit a query $(A_{t},b_{t})$ to the cloud with a probability $p_{t}$ that increases with the volume radius of the set $\mathcal{C}_{t}$, i.e., with $|\mathcal{C}_{t}|^{1/r_{t}}$ where
$|\mathcal{C}_{t}|=\frac{\pi^{r_{t}/2}}{\Gamma\big{(}r_{t}/2+1\big{)}}\cdot|\Lambda_{t}|^{1/2}\cdot(-\log\lambda_{t})^{r_{t}/2},$ (22)
with $\Gamma(\cdot)$ being the Gamma function. In this way, cloud-to-edge feedback is more likely to be requested at time instants $t$ characterized by a larger uncertainty, hence making cloud feedback more valuable for calibration. Conversely, when the HPD set is small enough, we reduce the probability of requesting feedback, thereby saving cloud computing resources.
@@
$p_{t}=\sigma\left(\frac{\log|\mathcal{C}_{t}|}{r_{t}}-\theta\right),$ (23)
which acts as a soft threshold on the log-volume radius. Lower values of the threshold $\theta$ in (23) increase the probability of cloud-to-edge feedback, yielding a smaller set $\mathcal{C}_{t}$ at the cost of higher usage of cloud resources.
@@
10:         $\lambda_{t+1}=\lambda_{t}-\frac{\gamma}{p_{t}}(\text{err}_{t}-\alpha)\text{obs}_{t}$triangleright update $\lambda_{t}$
11:     else
12:         $\lambda_{t+1}=\lambda_{t}$
13:     end if
14:end for